# MI-HGNN: Morphology-Informed Heterogeneous Graph Neural Network for Legged Robot Contact Perception

### v2

@@ II RELATED WORK
Learning-based contact perception: Learning-based contact perception involves learning techniques for robots to learn contact events through data-driven models. These models are trained on datasets of interactions between the robot and its environment, allowing the robot to adapt to complex and dynamic conditions. Learning-based contact perception is beneficial in scenarios where model-based approaches struggle, such as irregular terrains or unpredictable environments, as it allows the robot to improve the estimation of GRFs [19, 18] or contact state [17]. Learning-based methods face challenges due to their need for large labeled datasets, which are costly to collect in dynamic environments. They also function as “black-box” models, making them hard to interpret compared to physically grounded approaches. Additionally, they may ignore physical constraints, leading to unrealistic or unsafe behavior.
Geometric deep learning for robots: Instead of relying solely on the data to learn a model, geometric deep learning incorporates physical, geometric, or morphological constraints into the learning problem to reduce the search space and make the most of a small sample [23] [24]. Huang et al. propose “Shared Modular Policies”, a geometric framework where each segment in a variety of kinematically-diverse robots use a single policy for multi-task learning [25]. Ordonez et al. propose a theoretical framework for identifying discrete morphological symmetries (DMSs) of a dynamical system as a symmetry group $\mathcal{G}$, which can be exploited for augmenting robotic datasets or constraining learning models [26]. A subset of geometric deep learning includes graph neural networks (GNNs), which can model relational, temporal, or spatial relationships between nodes and exploit these relationships to learn robust policies. GNNs have been used to model robot kinematic chains [27] [28] [29], conduct multi-robot communication [30], learn data embeddings [31], and infer physical properties about particles [32] and simulated robots [33]. Inspired by the morphological symmetries found in legged robots, our MI-HGNN imposes shared weights between morphologically-identical kinematic branches to improve sample efficiency.
@@ III-A Legged Robot Dynamics
Legged robots such as quadrupeds and humanoids are typically treated as floating-base multi-body systems and are described by $n_{b}$ unactuated base coordinates $(\boldsymbol{q}_{b})$ and $n_{j}$ actuated joint coordinates $(\boldsymbol{q}_{j})$; as $\boldsymbol{q}=\begin{bmatrix}\boldsymbol{q}_{b},\boldsymbol{q}_{j}\end{bmatrix}^{T}$. In these generalized coordinates, $\boldsymbol{q}_{b}\in SE(3)$ represents the translation and orientation of the floating base $\mathcal{B}$ w.r.t. the inertial frame $\mathcal{I}$. Each value in $\boldsymbol{q}_{j}\in\mathbb{R}^{n_{j}}$, as $\boldsymbol{q}_{j,i}$, corresponds to the 1 DoF of the $i$th joint $\mathcal{J}_{i}$. The generalized velocity, acceleration, and torques of the system are $\dot{\boldsymbol{q}}$, $\ddot{\boldsymbol{q}}$, and $\boldsymbol{\tau}$, respectively.
The dynamics for this floating-base system are:
@@
where $\boldsymbol{M}(\boldsymbol{q})$ is the mass matrix, $\boldsymbol{C}(\boldsymbol{q},\dot{\boldsymbol{q}})$ are the coriolis and centrifugal terms, $\boldsymbol{g}(\boldsymbol{q})$ are the gravity terms, $\boldsymbol{S}^{T}$ is the selection matrix of actuated joints, $\boldsymbol{F}_{\textit{ext}}$ are the external forces, and ${\boldsymbol{J}_{\textit{ext}}({\boldsymbol{q}})}$ are the external force contact Jacobians [16] [20].
@@ III-B Model-Based Contact Force
This floating-base model has been used in many previous works for contact state estimation or GRF calculation on a quadruped with slightly modified assumptions [15] [16] [21]. Based on these formulations, each of the legs on a legged robot share a similar underlying function for contact force calculation. This relationship inspires us to share our MI-HGNN model parameters amongst the system’s legs to learn a generalizable dynamics equation similar to Eq. 3.
@@ IV-A Robot Morphology as a Heterogeneous Graph
We construct this heterogeneous graph directly from a robot’s kinematic structure; forming the node set $\mathcal{V}$ from the joints and the edge set $\mathcal{E}$ from the links. It is worth noting that among the joints we used to form the node set, the robot’s base body and feet are fixed joints whereas others are all revolute joints with an actuator. This is also one factor that inspires us to design a heterogeneous graph instead of a homogeneous graph in [25]. After construction, the connectivity of this heterogeneous graph is in accord with the specific robot morphology, which makes our neural network and learning algorithm morphology-aware.
@@ IV-B Sensor Fusion as Message Passing
where $\boldsymbol{v}_{i}$ is the embeddings of node $v_{i}$, $N_{A}(i)$ is the set of neighbor nodes connected to node $v_{i}$ of type $A$, and $\boldsymbol{W}$ is a learned matrix that depends on the current edge type $\mathcal{R}=\langle A,\phi(v_{i})\rangle$. We use a form of the graph operator from [35] that is modified for use with heterogeneous message passing.
@@ V-B Ground Reaction Force Estimation
Implementation: The input data for the MLP is a history of 150 samples that are flattened into one long array $\boldsymbol{x}=\mathbb{F}(\{\boldsymbol{z}_{i}\}_{i=0}^{150})\in\mathbb{R}^{6300}$, $\boldsymbol{z}$=$[\boldsymbol{q},\dot{\boldsymbol{q}},\boldsymbol{\tau},\boldsymbol{a}_{b},\boldsymbol{\omega}_{b}]\in\mathbb{R}^{42}$. For computing model-based results, the input $\boldsymbol{x}$ are transformed into the generalized coordinates: velocities, accelerations, and torques ($\boldsymbol{s},\dot{\boldsymbol{s}},\ddot{\boldsymbol{s}},\boldsymbol{\tau}_{s}$); with the base linear velocity, base angular acceleration, and joint acceleration estimated by taking derivatives of available measurements. Similar to the contact detection experiment, we re-organize the input data $\boldsymbol{x}$ into a graph structure for our MI-HGNN based on the node index and type, with minor modifications. Joint torques are also included in the joint node input, and there is no input to the foot node. Both MLP and MI-HGNN are trained with a learning rate of $10^{-4}$ and a batch size of 30, using the MSE loss as described in Section IV-C. The MLP uses a hidden size of 200 with 10 layers, and MI-HGNN uses a hidden size of 128 with 8 layers, leading to similar parameter sizes as listed in Table I.

### v3

@@ II RELATED WORK
Learning-based contact perception: Learning-based contact perception involves learning techniques for robots to learn contact events through data-driven models. These models are trained on datasets of interactions between the robot and its environment, allowing the robot to adapt to complex and dynamic conditions. Learning-based contact perception is beneficial in scenarios where model-based approaches struggle, such as irregular terrains or unpredictable environments, as it allows the robot to improve the estimation of GRFs [19, 18, 23] or contact state [17]. Learning-based methods face challenges due to their need for large labeled datasets, which are costly to collect in dynamic environments. They also function as “black-box” models, making them hard to interpret compared to physically grounded approaches. Additionally, they may ignore physical constraints, leading to unrealistic or unsafe behavior.
Geometric deep learning for robots: Instead of relying solely on the data to learn a model, geometric deep learning incorporates physical, geometric, or morphological constraints into the learning problem to make the most of a small sample [24, 25, 26]. Ordonez et al. propose a theoretical framework for identifying discrete morphological symmetries (DMSs) of a dynamical system as a symmetry group $\mathcal{G}$, which can be exploited for augmenting robotic datasets or constraining learning models [27]. A subset of geometric deep learning models, graph neural networks (GNNs), can model relational, temporal, or spatial relationships between nodes and exploit these relationships to learn complex functions encountered in robotics [28, 29, 30, 31]. Similar to our method, WAGNN [32] and Sanchez-Gonzalez et al. [33] both construct GNN architectures from the robot morphology, but employ homogeneous graphs with no ability to extract distinct features for different node types. NerveNet [34] also leverages a GNN derived from the kinematic tree for reinforcement learning, but differs from our method by casting both links and joints into nodes, resulting in a less compact graph compared to ours. Inspired by the morphological symmetries found in legged robots, our MI-HGNN imposes shared weights between morphologically-identical kinematic branches to improve model generalization and sample efficiency.
@@ III-A Legged Robot Dynamics
Legged robots such as quadrupeds and humanoids are typically treated as floating-base multi-body systems and are described by $n_{b}$ unactuated base coordinates $(\boldsymbol{q}_{b})$ and $n_{j}$ actuated joint coordinates $(\boldsymbol{q}_{j})$; as $\boldsymbol{q}=\begin{bmatrix}\boldsymbol{q}_{b},\boldsymbol{q}_{j}\end{bmatrix}^{T}$. In these generalized coordinates, $\boldsymbol{q}_{b}\in SE(3)$ represents the translation and orientation of the floating base $\mathcal{B}$ w.r.t. the inertial frame $\mathcal{I}$. Each value in $\boldsymbol{q}_{j}\in\mathbb{R}^{n_{j}}$, as $\boldsymbol{q}_{j,i}$, corresponds to the 1 DoF of the $i$th joint $\mathcal{J}_{i}$. The generalized velocity, acceleration, and torques of the system are $\dot{\boldsymbol{q}}$, $\ddot{\boldsymbol{q}}$, and $\boldsymbol{\tau}$, respectively. The dynamics for this floating-base system are:
@@
where $\boldsymbol{M}(\boldsymbol{q})$ is the mass matrix, $\boldsymbol{C}(\boldsymbol{q},\dot{\boldsymbol{q}})$ are the coriolis and centrifugal terms, $\boldsymbol{g}(\boldsymbol{q})$ are the gravity terms, $\boldsymbol{S}^{T}$ is the selection matrix of actuated joints, $\boldsymbol{F}_{\textit{ext}}$ are the external forces, and ${\boldsymbol{J}_{\textit{ext}}({\boldsymbol{q}})}$ are the external force contact Jacobians [16, 20].
@@ III-B Model-Based Contact Force
This floating-base model has been used in many previous works for contact state estimation or GRF calculation on a quadruped with slightly modified assumptions [15, 16, 21]. Based on these formulations, each of the legs on a legged robot share a similar underlying function for contact force calculation. This relationship inspires us to share our MI-HGNN model parameters amongst the system’s legs to learn a generalizable dynamics equation similar to Eq. 3.
@@ IV-A Robot Morphology as a Heterogeneous Graph
We construct this heterogeneous graph directly from a robot’s kinematic structure; forming the node set $\mathcal{V}$ from the joints and the edge set $\mathcal{E}$ from the links. We note that the robot’s base body and feet are fixed joints whereas others are all revolute joints with an actuator, which is one factor that inspires us to design a heterogeneous graph. After construction, the connectivity of this heterogeneous graph is in accord with the specific robot morphology, which makes our network and learning algorithm morphology-aware.
@@ IV-B Sensor Fusion as Message Passing
where $\boldsymbol{v}_{i}$ is the embeddings of node $v_{i}$, $N_{A}(i)$ is the set of neighbor nodes connected to node $v_{i}$ of type $A$, and $\boldsymbol{W}$ is a learned matrix that depends on the current edge type $\mathcal{R}=\langle A,\phi(v_{i})\rangle$. We use a form of the graph operator from [36] that is modified for use with heterogeneous message passing. Our heterogeneous encoders $\boldsymbol{E}_{A}$ and edge-informed aggregation strategy can enable more effective multi-modal sensor fusion than their homogeneous counterparts used in [34], as node features are computed with explicit modality awareness.
@@ V-B Ground Reaction Force Estimation
Implementation: The input data for the MLP is a history of 150 samples that are flattened into one long array $\boldsymbol{x}=\mathbb{F}(\{\boldsymbol{z}_{i}\}_{i=0}^{150})\in\mathbb{R}^{6300}$, $\boldsymbol{z}$=$[\boldsymbol{q},\dot{\boldsymbol{q}},\boldsymbol{\tau},\boldsymbol{a}_{b},\boldsymbol{\omega}_{b}]\in\mathbb{R}^{42}$. For computing model-based results, the input $\boldsymbol{x}$ are transformed into the generalized coordinates: velocities, accelerations, and torques ($\boldsymbol{s},\dot{\boldsymbol{s}},\ddot{\boldsymbol{s}},\boldsymbol{\tau}_{s}$); with the base linear velocity, base angular acceleration, and joint acceleration estimated by taking derivatives of available measurements. Similar to the contact detection experiment, we re-organize the input data $\boldsymbol{x}$ into a graph structure for our MI-HGNN, with minor modifications to append joint torques to the joint node input and exclude foot inputs. Both MLP and MI-HGNN are trained with a learning rate of $10^{-4}$ and a batch size of 30, using the MSE loss as described in Section IV-C. The MLP uses a hidden size of 200 with 10 layers, and the MI-HGNN uses a hidden size of 128 with 8 layers, leading to similar parameter sizes as listed in Table I.